OpenStudy (anonymous):

find the axis of symmetry for y=3x^2+5x-7 a)x=5 b)x=-5/3 c)x=-7 d)x=-5/6

OpenStudy (loser66):

What is the x coordinate of the vertex?

OpenStudy (anonymous):

it dosnt give me one

OpenStudy (loser66):

sure, you have to find it out

OpenStudy (loser66):

by formula x = -b/2a

OpenStudy (anonymous):

huh im so sorry im wakeful at this lol

OpenStudy (anonymous):

awful*

OpenStudy (anonymous):

\[y =3x^2+5x -7\] do you know how to do the completing the square? Look at the attachment, it briefly shows you how to do it. So If I complete the square given in the question then it will be \[y = 3(x + \frac{ 5 }{ 3 })^2-\frac{ 25 }{ 9 } - 7\] the x you see in this equation has to be to which if you add 5/3 will give you 0 Sp bring 5/3 to the other side \[x = -\frac{ 5 }{ 3 }\] back to the original equation so when x is -5/3 your y will be: \[y = -\frac{ 88 }{ 9 }\] the x is the line of symmetry